# Robust Decision Transformer: Tackling Data Corruption in Offline RL via Sequence Modeling

### v1

@@ 3.2.2 Gaussian Weighted Learning
In DT, actions serve as the most crucial supervised information, acting as the labels. Consequently, erroneous actions can directly influence the model through backpropagation. In RDT, we predict both actions and rewards, with rewards helping to reduce overfitting to corrupted action labels. To further mitigate the impact of corrupt data, we aim to reduce the influence of unconfident action and reward labels that could erroneously guide policy learning.
To identify uncertain labels, we adopt a simple but effective method: we use the value of the sample-wise loss to adjust the weight for the DT loss. The underlying insight is that a corrupted label would typically lead to a larger loss. To softly reduce the effect of potentially corrupted labels, we use the Gaussian weight, i.e., a weight that decays exponentially in accordance with the sample’s loss. This is mathematically represented as:
@@
$\delta_{r_{t}}=\text{no\_grad}(\|\pi_{\theta}(\hat{\tau}_{t-K+1:t-1}^{(i)},\hat{R}_{t}^{(i)},\hat{s}_{t}^{(i)},\hat{a}_{t}^{(i)})-\hat{r}_{t}^{(i)}\|_{2}),$
In this equation, $\delta_{a_{t}}$ and $\delta_{r_{t}}$ represent the detached prediction errors at step $t$. The variables $\beta_{a}\geq 0,\beta_{r}\geq 0$ act as the temperature coefficients, providing flexibility to control the "blurring" effect of the Gaussian weights. With a larger value, more samples would be down-weighted. The loss function of RDT is expressed as follows:
@@ 3.2.3 Iterative Data Correction
While Gaussian weighted learning has significantly reduced the impact of corrupted labels, these erroneous data can still influence the policy as they still serve as inputs for policy learning. We propose iteratively correcting corrupted data in the dataset using the model’s predictions to bring the data closer to their true values in the next iteration. This method can further minimize the detrimental effects of corruption and can be implemented to correct reward-to-gos, states, and actions in the sequence modeling framework. In RDT, it is straightforward to correct the actions and rewards in the dataset, and recalculate the reward-to-gos using corrected rewards. Therefore, our implementation focuses on correcting states and reward-to-gos in the datasets, leaving better data correction methods for states for future work.
@@
Initially, we store the distribution information of prediction error $\delta$ in Eq.4 throughout the learning phase to preserve the mean $\mu_{\delta}$ and variance $\sigma^{2}_{\delta}$ of actions and rewards, updating them with every batch of samples. Our hypothesis is that the prediction error $\delta$ between predicted and clean label actions should exhibit consistency after sufficient training. Therefore, if erroneous label actions are encountered, $\delta$ will deviate from the mean $\mu_{\delta}$, behaving like outliers. This deviation essentially enables us to detect and correct the corrupted action using the distributional information of $\delta$.
@@
Taking actions as an example, to detect corrupted actions, we calculate the $z$-score, denoted by $z^{(i)}=\frac{\delta^{(i)}-\mu_{\delta}}{\sigma_{\delta}}$, for each sampled action $\hat{a}^{(i)}$ and the current policy. If the condition $z^{(i)}>\zeta\cdot\sigma_{\delta}$ is met for any given action $\hat{a}^{(i)}$, we infer that the action $\hat{a}^{(i)}$ has been corrupted. We then permanently replace $\hat{a}^{(i)}$ in the dataset with the predicted action. Correcting rewards follows a similar process, with the additional step of recalculating the reward-to-gos. The hyperparameter $\zeta$ determines the detection thresholds: a smaller $\zeta$ will classify more samples as corrupted data, while a larger $\zeta$ results in fewer modifications to the dataset.
@@ 4 Experiments
In this section, we conduct a comprehensive empirical assessment of RDT to address the following questions: 1) How does RDT perform under several data corruption scenarios during the training phase? 2) Is RDT robust to different data corruption rates and scales? 3) What is the impact of each component of RDT on its overall performance? 4) How does RDT perform when faced with observation perturbations during the testing phase?
@@ 4.1 Experimental Setups
We evaluate RDT across a variety of tasks, such as MoJoCo ("medium-replay-v2" datasets), KitChen, and Adroit (Fu et al.,, 2020). Two types of data corruption during the training phase are simulated: random corruption and adversarial corruption, attacking states, actions, and rewards, respectively. The corruption rate is set to $0.3$ and the corruption scale is $\epsilon=1.0$ for the main results. Data corruption is introduced in Section 2.3 and more details can be found in Section C.1. Additionally, we find that the data corruption problem can exacerbate when the data is limited. To simulate this scenario, we down-sample on both MuJoCo and Adroit tasks, referred to as MuJoCo (10%) and Adroit (1%), Specifically, we randomly select 10% (and 1%) of the trajectories from MuJoCo (and Adroit) tasks and conduct data corruption on the sampled data. We do not down-sample the Kitchen dataset because it already has a limited dataset size. All considered tasks have a similar dataset size of about 20,000 transitions, except for the original MuJoCo task, which includes a larger dataset.
We compare RDT with several SOTA offline RL algorithms and robust methods, namely BC, RBC (Sasaki and Yamashina,, 2020), DeFog (Hu et al.,, 2023), CQL (Kumar et al.,, 2020), UWMSG (Ye et al., 2024b, ), RIQL (Yang et al., 2024b, ), and DT (Chen et al.,, 2021). BC and RBC employ behavior cloning loss within an MLP-based model for policy learning, while DeFog and DT utilize a Transformer architecture. We implement RBC with our Gaussian-weighted learning as an instance of  (Sasaki and Yamashina,, 2020). In Appendix D.1, we investigate the robustness of DT across various critical parameters. Drawing from these findings on DT’s robustness, we establish a default implementation of DT, DeFog, RDT with a sequence length of 20 and a block number of 3. To ensure the validity of the findings, each experiment is repeated using 4 different random seeds.
Table 1: Average normalized performance under random data corruption.
Task Attack Element BC RBC DeFog CQL UWMSG RIQL DT RDT (ours) MuJoCo state 24.1 31.5 19.5 22.0 5.7 22.7 45.5 49.1 action 26.9 29.6 49.7 56.9 52.3 57.8 52.4 54.1 reward 28.1 29.1 42.9 58.6 51.3 60.4 52.7 54.8 MuJoCo (10%) state 9.5 8.0 16.2 4.5 6.8 9.7 21.4 23.1 action 9.0 7.4 15.4 10.3 19.3 12.8 19.4 25.9 reward 10.0 11.0 16.2 12.7 9.4 18.7 27.5 31.6 KitChen state 39.9 29.8 20.8 5.1 0.1 22.5 31.2 43.1 action 37.9 37.3 13.7 3.6 0.0 30.6 28.8 40.4 reward 45.3 50.3 19.7 0.1 0.0 43.0 49.2 52.6 Adroit (1%) state 54.5 74.2 72.9 -0.1 -0.1 23.1 86.9 96.4 action 35.9 71.8 83.9 -0.1 -0.1 12.6 76.8 99.1 reward 34.7 88.6 88.8 -0.1 -0.1 22.7 84.7 100.9 Average 29.6 39.0 38.3 14.5 12.1 28.1 48.0 55.9
@@ 4.2 Evaluation under Various Data Corruption
Random Data Corruption.
@@
To address the first question, we evaluate the RDT and baselines on random data corruption across a diverse set of task groups. Each group consists of three datasets, and the mean normalized score was calculated to represent the final outcomes. As shown in Table 1, RDT demonstrates superior performance in handling data corruption, achieving the highest score in 10 out of 12 settings, particularly in the challenging Kitchen and Adroit tasks. Notably, across all task groups, RDT outperforms DT, underscoring its effectiveness in reducing DT’s sensitivity to data corruption across all elements. Moreover, RBC significantly outperforms BC, further highlighting the effectiveness of Gaussian weighted learning. However, prior offline RL methods, such as RIQL and UWMSG, fail to yield satisfactory results and even underperform BC overall, despite showing enhanced performance on the full MuJoCo dataset. This indicates that temporal difference methods are dependent on large-scale datasets.
Figure 3: Results under mixed random data corruption.
Mixed Random Data Corruption.
[…]
Adversarial Data Corruption.
We further extend the analysis to examine the robustness of the RDT under adversarial data corruption. We employ the same hyperparameters for RDT as those utilized in the random data corruption scenarios, demonstrating RDT’s advantage of not requiring meticulous hyperparameter tuning. As evidenced in Table 2, RDT consistently showcases a robust performance by attaining the highest scores in 8 out of 9 settings and sustaining the top average performance. Notably, RDT stands out for enhancing the average score by $28.9\%$ compared to DT in the adversarial action corruption scenario. Intriguingly, temporal-difference methods like CQL, UWMSG, and RIQL perform notably worse than BC and sequence modeling methods such as DT and DeFog, highlighting the promise of sequence modeling approaches. We provide detailed results for each task in Appendix D for comprehensive analysis.
Table 2: Average normalized performance under adversarial data corruption.
@@
Task Attack Element BC RBC DeFog CQL UWMSG RIQL DT RDT (ours) MuJoCo (10%) state 10.8 10.9 8.7 4.5 13.4 11.0 22.8 22.6 action 5.6 6.4 10.4 11.4 18.1 7.1 12.9 17.5 reward 10.0 10.8 13.8 11.7 7.8 18.3 27.7 29.1 KitChen state 33.5 37.2 20.8 1.8 0.0 28.5 36.4 42.3 action 5.0 7.8 13.7 0.7 5.1 4.4 8.1 33.7 reward 45.3 43.4 19.7 0.0 0.0 38.8 49.8 54.0 Adroit (1%) state 66.6 70.5 80.5 -0.1 -0.1 37.0 90.4 96.8 action 19.3 24.3 54.2 -0.1 -0.1 4.4 35.2 90.2 reward 34.7 84.0 82.9 -0.1 -0.1 26.4 87.9 91.4 Average 25.6 32.8 33.9 3.3 4.9 19.5 41.2 53.1
4.3 Varying Corruption Rates and Scales
In the above experiments, we evaluate the effectiveness of RDT under a data corruption rate of $0.3$ and a scale of $1.0$. We further examine the robustness of RDT under different corruption rates from $\{0.0,0.1,0.3,0.5\}$ and scales from $\{0.0,1.0,2.0\}$. As illustrated in Figure 4, RDT consistently delivers superior performance compared to other baselines across different corruption rates and scales. However, we did observe a decline in RDT’s performance when it encountered state corruption with high rates. This decline can be attributed to the distortion of the state’s distribution caused by the high corruption rates, resulting in a significant deviation from the clean test environment to the corrupted datasets. Enhancing the robustness of RDT against state corruption, especially under high corruption rates and scales, holds potential for future advancements.
Figure 4: Results under various data corruption rates (a) and scales (b) on "walker2d-medium-replay".

### v3

@@ 3.2.2 Gaussian Weighted Learning
In addition to mitigating the negative impact of erroneous inputs, it is also essential to address the influence of corrupted labels. In DT, actions serve as the most crucial supervised signals, acting as the labels. Therefore, erroneous actions can directly influence the model through backpropagation. In RDT, we predict both actions and rewards, with rewards helping to reduce overfitting to corrupted action labels. To further diminish the impact of corrupt data, we focus on minimizing the influence of unconfident action and reward labels that could misguide policy learning.
To identify uncertain labels, we adopt a simple but effective method: we use the value of the sample-wise loss to adjust the weight for the DT loss. The underlying insight is that a corrupted label would typically lead to a larger loss (see Figure 3(b)). To softly reduce the effect of potentially corrupted labels, we use the Gaussian weight, i.e., a weight that decays exponentially in accordance with the sample’s loss. This is mathematically formulated as:
@@
$\delta_{r_{t}}=\text{no\_grad}(\|\pi_{\theta}(\hat{\tau}_{t-K+1:t-1}^{(i)},\hat{R}_{t}^{(i)},\hat{s}_{t}^{(i)},\hat{a}_{t}^{(i)})-\hat{r}_{t}^{(i)}\|_{2}).$
In Eq.5, $\delta_{a_{t}}$ and $\delta_{r_{t}}$ represent prediction errors at step $t$ with detached gradients. The variables $\beta_{a}\geq 0,\beta_{r}\geq 0$ act as the temperature coefficients, providing flexibility to control the “blurring” effect of the Gaussian weights. Different tasks exhibit preferences for different temperature coefficients, as demonstrated in Appendix D.3. With a larger value, more samples would be down-weighted. The loss function of RDT is expressed as follows:
@@ 3.2.3 Iterative Data Correction
While the first two techniques have significantly reduced the impact of corrupted data, these erroneous data can still affect the policy as they remain in the datasets. We propose iteratively correcting corrupted data in the dataset using the model’s predictions to bring the data closer to their true values in the next iteration. This method can further minimize the detrimental effects of corruption and can be implemented to correct reward-to-go, state, and action elements. In RDT, it is straightforward to correct the actions and rewards in the dataset and recalculate the reward-to-gos using corrected rewards. Therefore, our implementation focuses on correcting actions and reward-to-gos in the datasets, leaving better data correction methods for states for future work.
@@
Initially, we store the distribution information of prediction error $\delta$ in Eq.5 throughout the learning phase to preserve the mean $\mu_{\delta}$ and variance $\sigma^{2}_{\delta}$ of actions and rewards, updating them with every batch of samples. The hypothesis is that the prediction error $\delta$ between predicted and clean labels should exhibit consistency after sufficient training. Therefore, if erroneous label actions are encountered, $\delta$ will deviate from the mean $\mu_{\delta}$, behaving like outliers. This deviation essentially enables us to detect and correct the corrupted data using the information of $\delta$.
@@
Taking actions as an example, to detect corrupted actions, we calculate the $z$-score, denoted by $z^{(i)}=\frac{\delta^{(i)}-\mu_{\delta}}{\sigma_{\delta}}$, for each sampled action $\hat{a}^{(i)}$. If the condition $z^{(i)}>\zeta\cdot\sigma_{\delta}$ is met for any given action $\hat{a}^{(i)}$, we infer that the action $\hat{a}^{(i)}$ has been corrupted. We then permanently replace $\hat{a}^{(i)}$ in the dataset with the predicted action. This helps eliminate corrupted actions from the dataset (see Figure 3(c)). Correcting rewards follows a similar process, with the additional step of recalculating the reward-to-gos. The hyperparameter $\zeta$ determines the detection thresholds: a smaller $\zeta$ will classify more samples as corrupted data, while a larger $\zeta$ results in fewer modifications to the dataset. Empirically, we can achieve performance improvement by setting $\zeta$ to approximately 5. More implementation details of RDT can be found in Appendix C.3.
Table 1: Detailed comparative results under random data corruption.
Attack Task BC RBC DeFog CQL UWMSG RIQL DT RDT State halfcheetah (10%) 1.9$\pm$ 0.2 1.7$\pm$ 0.5 6.7$\pm$ 0.5 9.7$\pm$ 0.8 3.9$\pm$ 0.5 4.4$\pm$ 0.9 6.3$\pm$ 0.4 8.3$\pm$ 1.5 hopper (10%) 16.4$\pm$ 3.7 13.0$\pm$ 2.3 23.1$\pm$ 5.5 2.2$\pm$ 1.5 16.3$\pm$ 7.4 15.5$\pm$ 5.4 36.1$\pm$ 7.6 40.8$\pm$ 3.5 walker2d (10%) 7.5$\pm$ 0.9 10.3$\pm$ 1.4 9.0$\pm$ 3.3 1.2$\pm$ 1.8 -0.4$\pm$ 0.2 9.2$\pm$ 4.4 18.0$\pm$ 2.5 20.3$\pm$ 2.8 kitchen-complete 15.9$\pm$ 5.2 21.1$\pm$ 3.1 37.1$\pm$ 13.4 12.9$\pm$ 4.3 0.0$\pm$ 0.0 37.5$\pm$ 6.4 37.0$\pm$ 6.2 52.8$\pm$ 1.8 kitchen-partial 17.1$\pm$ 2.6 20.2$\pm$ 2.8 9.8$\pm$ 6.6 0.0$\pm$ 0.0 0.0$\pm$ 0.0 25.9$\pm$ 3.4 31.0$\pm$ 8.1 36.8$\pm$ 5.8 kitchen-mixed 17.4$\pm$ 4.9 29.2$\pm$ 5.6 10.6$\pm$ 5.0 0.0$\pm$ 0.0 0.0$\pm$ 0.0 21.6$\pm$ 3.7 31.8$\pm$ 3.4 41.8$\pm$ 4.3 door (1%) 75.7$\pm$ 11.7 73.3$\pm$ 9.2 101.3$\pm$ 1.5 -0.3$\pm$ 0.0 -0.2$\pm$ 0.2 39.0$\pm$ 16.4 94.6$\pm$ 4.2 102.8$\pm$ 2.4 hammer (1%) 90.5$\pm$ 7.8 90.6$\pm$ 3.5 93.0$\pm$ 15.8 0.2$\pm$ 0.0 -0.0$\pm$ 0.1 70.0$\pm$ 12.6 97.8$\pm$ 12.3 113.8$\pm$ 1.6 relocate (1%) 7.2$\pm$ 4.5 12.2$\pm$ 6.8 14.7$\pm$ 3.7 -0.2$\pm$ 0.0 -0.3$\pm$ 0.0 5.2$\pm$ 5.0 61.6$\pm$ 5.6 65.0$\pm$ 6.2 Average 27.2 30.2 33.9 2.8 2.2 25.4 46.0 53.6 Action halfcheetah (10%) 0.8$\pm$ 0.3 0.7$\pm$ 0.1 10.3$\pm$ 1.8 19.8$\pm$ 5.1 8.0$\pm$ 0.6 2.3$\pm$ 0.6 3.6$\pm$ 0.3 20.4$\pm$ 1.7 hopper (10%) 18.3$\pm$ 2.4 17.7$\pm$ 2.3 32.8$\pm$ 3.9 1.8$\pm$ 0.0 33.0$\pm$ 4.3 26.3$\pm$ 2.8 32.1$\pm$ 6.0 37.6$\pm$ 5.7 walker2d (10%) 5.1$\pm$ 1.4 6.1$\pm$ 3.5 13.9$\pm$ 9.6 3.1$\pm$ 2.4 6.4$\pm$ 0.5 9.8$\pm$ 1.9 22.2$\pm$ 4.0 30.4$\pm$ 2.5 kitchen-complete 3.5$\pm$ 2.0 4.6$\pm$ 1.2 23.0$\pm$ 14.5 12.6$\pm$ 4.2 0.0$\pm$ 0.0 20.1$\pm$ 6.1 12.6$\pm$ 5.2 44.0$\pm$ 3.7 kitchen-partial 29.9$\pm$ 3.0 34.0$\pm$ 3.5 5.9$\pm$ 2.5 5.4$\pm$ 6.3 0.0$\pm$ 0.0 33.4$\pm$ 5.6 26.2$\pm$ 0.8 40.8$\pm$ 3.7 kitchen-mixed 34.9$\pm$ 3.8 39.5$\pm$ 4.5 17.0$\pm$ 7.7 0.0$\pm$ 0.0 0.0$\pm$ 0.0 41.0$\pm$ 10.4 30.4$\pm$ 2.2 44.4$\pm$ 3.7 door (1%) 45.0$\pm$ 15.4 61.4$\pm$ 9.1 100.7$\pm$ 3.1 -0.3$\pm$ 0.1 -0.2$\pm$ 0.1 29.6$\pm$ 29.1 84.3$\pm$ 5.5 103.3$\pm$ 1.0 hammer (1%) 76.4$\pm$ 15.2 67.8$\pm$ 6.9 99.7$\pm$ 17.8 0.2$\pm$ 0.0 0.1$\pm$ 0.1 68.7$\pm$ 29.4 72.0$\pm$ 5.8 116.3$\pm$ 0.8 relocate (1%) 26.0$\pm$ 10.4 44.3$\pm$ 7.4 44.9$\pm$ 2.6 -0.2$\pm$ 0.1 -0.3$\pm$ 0.0 20.4$\pm$ 9.6 63.2$\pm$ 4.4 69.9$\pm$ 9.7 Average 26.6 30.7 38.7 4.7 5.2 28.0 38.5 56.3 Reward halfcheetah (10%) 2.4$\pm$ 0.2 2.9$\pm$ 0.3 14.3$\pm$ 3.6 30.4$\pm$ 1.9 2.2$\pm$ 0.7 6.1$\pm$ 1.2 9.3$\pm$ 0.9 21.7$\pm$ 4.0 hopper (10%) 19.7$\pm$ 2.8 19.3$\pm$ 2.3 22.8$\pm$ 4.2 1.8$\pm$ 0.0 18.2$\pm$ 2.8 40.2$\pm$ 2.6 36.0$\pm$ 7.4 42.8$\pm$ 3.4 walker2d (10%) 9.7$\pm$ 1.5 8.3$\pm$ 3.8 14.2$\pm$ 2.1 1.2$\pm$ 0.7 7.9$\pm$ 0.8 9.6$\pm$ 2.5 27.4$\pm$ 3.4 30.1$\pm$ 4.5 kitchen-complete 36.0$\pm$ 11.5 38.2$\pm$ 8.4 43.2$\pm$ 5.7 1.4$\pm$ 2.4 0.0$\pm$ 0.0 52.8$\pm$ 6.8 43.9$\pm$ 4.3 65.6$\pm$ 4.4 kitchen-partial 34.1$\pm$ 1.4 39.1$\pm$ 3.2 7.9$\pm$ 6.4 0.0$\pm$ 0.0 0.0$\pm$ 0.0 36.4$\pm$ 2.2 47.1$\pm$ 6.9 51.4$\pm$ 2.0 kitchen-mixed 38.9$\pm$ 1.4 47.1$\pm$ 1.9 14.9$\pm$ 7.2 0.0$\pm$ 0.0 0.0$\pm$ 0.0 49.8$\pm$ 4.3 42.8$\pm$ 1.9 47.8$\pm$ 4.3 door (1%) 76.0$\pm$ 5.9 75.0$\pm$ 3.2 102.3$\pm$ 2.7 -0.3$\pm$ 0.0 -0.2$\pm$ 0.0 56.1$\pm$ 9.5 99.0$\pm$ 2.3 103.7$\pm$ 0.8 hammer (1%) 97.1$\pm$ 8.3 99.0$\pm$ 5.1 101.8$\pm$ 10.2 0.2$\pm$ 0.0 0.4$\pm$ 0.4 52.2$\pm$ 22.7 80.7$\pm$ 11.2 123.8$\pm$ 2.2 relocate (1%) 36.1$\pm$ 8.6 32.2$\pm$ 6.7 53.1$\pm$ 3.1 -0.3$\pm$ 0.1 -0.3$\pm$ 0.0 9.1$\pm$ 5.3 71.8$\pm$ 7.7 84.9$\pm$ 1.6 Average 38.9 40.1 41.6 3.8 3.1 34.7 50.9 63.5 Average over all tasks 31.1 33.7 38.1 3.8 3.5 29.3 45.1 57.8
@@ 4 Experiments
In this section, we conduct a comprehensive empirical evaluation of RDT, focusing on the following key questions: 1) How does RDT perform under different data corruption scenarios? 2) Is RDT robust to observation perturbations during the testing phase? 3) What is the contribution of each component of RDT to its overall performance? 4) How do baselines compare to RDT when equipped with a transformer backbone?
@@ 4.1 Experimental Setups
We evaluate RDT using various offline RL benchmarks, such as MuJoCo, KitChen, and Adroit (Fu et al., 2020). Two types of data corruption during the training phase are simulated: random and adversarial corruption, attacking states, actions, and rewards. Specifically, the corruption rate is set to $30\%$, and the corruption scale is $\epsilon=1.0$ for the main results , similar to previous work (Yang et al., 2024b). Data corruption is introduced in Section 2, and more details can be found in Section C.1. We find that the data corruption problem can be exacerbated when the data is limited. To simulate such a scenario, we down-sample on both MuJoCo and Adroit tasks, referred to as MuJoCo (10%) and Adroit (1%), Specifically, we randomly select 10% (and 1%) of the trajectories from MuJoCo (and Adroit) tasks and conduct data corruption on the sampled data. We refrain from down-sampling the Kitchen dataset due to its already limited size. Detailed information about the dataset sizes can be found in Section C.4.
We compare RDT with several SOTA offline RL algorithms and corruption-robust methods, namely BC, RBC (Sasaki & Yamashina, 2020), DeFog (Hu et al., 2023), CQL (Kumar et al., 2020), UWMSG (Ye et al., 2024b), RIQL (Yang et al., 2024b), and DT (Chen et al., 2021). BC and RBC employ behaviour cloning loss within an MLP-based model for policy learning, while DeFog and DT utilize a Transformer architecture. We implement RBC with our Gaussian-weighted learning as an instance of  (Sasaki & Yamashina, 2020). In Appendix D.1, we investigate the robustness of DT across various critical parameters. Drawing from these findings on DT’s robustness, we establish a default implementation of DT, DeFog, RDT with a sequence length of 20 and a block number of 3. To ensure the validity of the findings, each experiment is repeated using 4 different random seeds, and we also report the standard variance over these random seeds.
@@ 4.2 Evaluation under Various Data Corruption
Results under Random Corruption.
@@
To address the first question, we first evaluate the RDT and baselines under the random data corruption scenario. The mean of normalized scores across different seeds is calculated as the evaluation criterion for each task. As shown in Table 1, RDT demonstrates superior performance in handling data corruption, achieving the highest score in 24 out of 27 settings. Notably, across all tasks, RDT consistently outperforms DT, with a significant overall improvement of 28.2%, underscoring its effectiveness in reducing DT’s sensitivity to data corruption. Moreover, RBC slightly outperforms BC, further highlighting the effectiveness of the Gaussian weighted learning approach. However, prior offline RL methods, such as RIQL and UWMSG, fail to yield satisfactory results and even underperform BC, indicating temporal difference methods’ weakness in the limited data domain. We further evaluate RDT under varying corruption scales and ratios in Section D.4, where RDT demonstrates superior robustness compared to other baselines.
Table 2: Results under adversarial data corruption. Each score is averaged across the task group.
@@
Attack Task BC RBC DeFog CQL UWMSG RIQL DT RDT State MuJoCo (10%) 9. 9 9. 4 12. 9 4. 2 9. 0 11. 0 21. 9 23. 6 KitChen 23. 4 28. 9 20. 0 2. 2 0. 0 40. 4 37. 9 49. 1 Adroit (1%) 60. 2 53. 6 75. 6 -0. 1 -0. 1 44. 5 85. 3 95. 4 Average 31. 2 30. 6 36. 2 2. 1 2. 9 32. 0 48. 4 56. 0 Action MuJoCo (10%) 4. 2 4. 3 10. 3 7. 3 16. 9 7. 1 13. 4 21. 6 KitChen 6. 2 11. 5 3. 9 0. 9 0. 0 8. 0 5. 4 34. 0 Adroit (1%) 8. 4 20. 7 51. 8 -0. 1 -0. 1 42. 4 47. 4 80. 3 Average 6. 3 12. 2 22. 0 2. 7 5. 6 19. 2 22. 1 45. 3 Reward MuJoCo (10%) 10. 6 10. 2 11. 5 11. 8 15. 2 18. 3 25. 2 31. 9 KitChen 36. 3 41. 5 20. 3 1. 1 9. 7 45. 9 48. 1 56. 0 Adroit (1%) 69. 7 68. 7 80. 8 -0. 1 0. 0 53. 6 90. 9 96. 5 Average 38. 9 40. 1 37. 6 4. 3 -0. 1 39. 2 54. 7 61. 5 Average over all tasks 25. 5 27. 7 31. 9 3. 0 4. 4 30. 1 41. 7 54. 3
Results under Adversarial Corruption.
We further extend the analysis to examine the robustness of the RDT under an adversarial data corruption scenario. As illustrated in Table 2, we calculate the mean score across the task group. RDT consistently demonstrates robust performance, achieving the highest average scores under various attacked elements. Notably, RDT improves the average score by 105% compared to DT in the adversarial action corruption scenario. Intriguingly, temporal-difference methods like CQL and UWMSG perform significantly worse than both BC and sequence modeling methods such as DT and DeFog, underscoring the potential of sequence modeling approaches. Detailed results for each task are provided in Appendix D for comprehensive analysis.
Results under Mixed Corruption.
To present a more challenging scenario for assessing the robustness of RDT, we conduct experiments under mixed data corruption settings. In this setting, all three elements (states, actions, and rewards) are corrupted at a rate of $30\%$ and a scale of $1.0$. As shown in Figure 4, RDT consistently outperforms other baselines across all tasks, highlighting its superior stability even when faced with simultaneous and diverse data corruptions. Notably, in the challenging Kitchen and Adroit tasks with mixed adversarial settings, RDT surpasses DT and RIQL by an impressive margin of approximately 100%.
Figure 4: Results under (a) mixed random corruption and (b) mixed adversarial corruption.
We investigate the robustness of RDT when deployed in perturbed environments after being trained on corrupted data, a challenging setting that includes both training-time and testing-time attacks. To address this, we evaluate RDT under two types of observation perturbations during the testing phase: Random and Action Diff, following prior works (Yang et al., 2022a; Zhang et al., 2020). Among these, Action Diff involves sampling multiple random noises to maximize the impact on the difference in policy output. The perturbation scale is used to control the extent of influence on the observation. Details of observation perturbations are provided in Section C.2.
The comparison results are presented in Figure 5, where all algorithms are trained on the offline dataset with mixed random corruption and evaluated under observation perturbation. These results demonstrate the superior robustness of RDT under the two types of observation perturbations, maintaining stability even at a high perturbation scale of 0.5. Notably, RORL (Yang et al., 2022a), a SOTA offline RL method designed to tackle testing-time observation perturbation, fails on these tasks due to data corruption. Additionally, DT and RIQL experience significant performance drops as the perturbation scale increases. Additional results comparing algorithms trained on offline datasets with corrupted states, actions, or rewards can be found in Section D.7.
(a) Random observation perturbation.
(b) Action Diff observation perturbation.
Figure 5: Performance under various observation perturbation scales during the testing phase. All the algorithms are trained under mixed random corruption during the training phase.